OpenStudy (anonymous):

Use the first derivative to find all critical points and use the second derivative to find all inflection points. Use a graph to identify each critical point as a local maximum, a local minimum, or neither. f(x)=x^5-25x^4+50. Enter answers in increasing order.

OpenStudy (anonymous):

you there?

OpenStudy (anonymous1):

Derivative: f'(x) = 5x^4-100x^3

OpenStudy (anonymous):

f'(x) = 5x^4 - 100x^3 = 0 at critical points 5x^3(x - 20) = 0 x = 0 and x = 20 are critical points

OpenStudy (anonymous1):

Critical points are the points where (1) the derivative is zero or (2) the derivative doesn't exist.

OpenStudy (anonymous):

and?

OpenStudy (anonymous):

i also need inflection point and whether the criticals are local maximum or minimum

OpenStudy (anonymous1):

The second derivative is: 20x^3-300x^2

OpenStudy (anonymous1):

20x^3-300x^2=0 x^2(20x-300)=0 So, the zeroes of the second derivative are: x=0 and x=300/20=15

OpenStudy (anonymous):

second derivative is 20x^3 - 300 x^2 x=0 gives value of 0 x = 20 gives positive value so this is a minimum

OpenStudy (anonymous):

still not aswring the full question

OpenStudy (anonymous):

at x = 0 we could have point of inflection or a maximum

OpenStudy (anonymous):

ok so could u please just put every answer together at one time?

OpenStudy (anonymous):

check f'(x) around region of x = 0 f'(-0.1) = 5(-0.1)^4 - 100(0.01)^3 is positive f'(0.01) is negative thus x = 0 gives a local maxm

OpenStudy (anonymous):

ok you have a local maximum at (0,50) and a local minimum at (20, -799950)

OpenStudy (anonymous):

thank u! but what about the inflection point. What is that?

OpenStudy (anonymous):

i'm not sure if there is one - let me check

OpenStudy (anonymous):

thnks

OpenStudy (anonymous):

no - if second derivative = 0 as it is at x = 0 it could be a point of inflection but the slopes around x = 0 show a maximum - i plotted the function using wolfram alpha which confirmed this there is only 1 maxm and 1 minimum

OpenStudy (anonymous):

so no?

OpenStudy (anonymous1):

You can verify that x = 0 is not an inflection point by checking the values of the second derivative around x = 0 20(0.01)^3 - 300 (0.01)^2 = -0.02998 20(-0.01)^3 - 300 (-0.01)^2 = -0.03002 Since around x = 0 the second derivative does not change sign (it is always negative) it is not an inflection point.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok, no inflection point

OpenStudy (anonymous1):

We still have to check whether there is an inflection point at x =15.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

u mean 20

OpenStudy (anonymous1):

I mean 15, because it is the other zero of the second derivative. Let's check the signs of the second derivative around x = 15. 20(14.998)^3 - 300(14.998)^2 = -8.99760016000073 20(15.002)^3 - 300(15.002)^2 = 9.002400159995886 It changes signs around x=15, therefore it is an inflection point.

OpenStudy (anonymous):

so x=15 is the inflection point?

OpenStudy (anonymous1):

(15, -506200) is the inflection point.

OpenStudy (anonymous):

which one is it? x=?

OpenStudy (anonymous):

15?

OpenStudy (anonymous1):

Correct.

OpenStudy (anonymous):

im going to go try all this and giv u feedback

OpenStudy (anonymous1):

You can confirm that with Wolfram|Alpha: http://www.wolframalpha.com/input/?i=inflection+points+of+x^5-25x^4%2B50

OpenStudy (anonymous):

k one sec thanks

OpenStudy (anonymous):

the critical points are 0 with local max and 20 with local min?

OpenStudy (anonymous1):

Correct.

OpenStudy (anonymous):

ok, thank you!

OpenStudy (anonymous1):

You're welcome.

OpenStudy (anonymous):

than you also anonymous - i've learned something new about points of inflection today

OpenStudy (anonymous1):

You're welcome, joeywhite.